Using the below parameters calculate what is the budget surplus/deficit first a)in a closed economy with a government and b)in an open economy with a government.
(Note = C=autonomous consumption, I=Investment, G is government expenditure, X is exports, t is tax, Z is imports, Y is income.

C = 100 + 0.4Y
I = 300
G= 200
X = 300
T = 0.2
Z = 0.4

Please show workings. The answer for a) is 23.53 surplus and the answer for b) is also 23.53 surplus. However i am unclear how they got to this answer. Please assist . Thank you
1
Expert's answer
2019-01-11T12:32:11-0500

a) In a closed economy with a government:

Y = C + I + G = 100 + 0.4Y + 300 + 200,

0.6Y = 600,

Y = 1000,

T = t*Y = 0.2*1000 = 200.

The budget surplus/deficit is:

T - G = 200 - 200 = 0.

b) In an open economy with a government.

Y = C + I + G + (X - z*Y) = 100 + 0.4Y + 300 + 200 + (300 - 0.4Y),

Y = 900,

T = t*Y = 0.2*900 = 180.

The budget surplus/deficit is:

T - G = 180 - 200 = -20, so there is a deficit of 20.